o How does lift change with airspeed if a constant AOA and altitude are held? (provide specific examples)
Answer #1

We know that lift is in quadrtic relation with the air speed.
if we keep constant the other parameter then
Lift (L)= K*(Airspeed(V))2
hence the lift would vary parabolically with the air speed.
For e.g. say air speed becomes half
L1 = K*(V/2)2 = (1/4)KV2
L1 = (1/4)L
hence on halving the air speed lift would reduce to (1/4)th